OpenStudy (roadjester):

Decide whether the given vector is a member of Span(S) for each given set S. If so, express it as a linear combination from S:

OpenStudy (roadjester):

\[\huge S=\left\{ \frac1 x , \frac 1 {x^2}, \frac1 {x^3}\right\}\subseteq F((0,\infty)); f(x)= {2{x^2}-7x-10\over {x^3}}\]

OpenStudy (roadjester):

\[\huge f(x)={2{x^2}-7x-10\over {x^3}}\]

OpenStudy (anonymous):

Basically, you have to ask yourself, can:\[\frac{2x^2-7x-10}{x^3}\]be written as a linear combination of 1/x, 1/x^2, and 1/x^3? in other words, can you find scalars a, b, and c, such that: \[\frac{2x^2-7x-10}{x^3}=a\left(\frac{1}{x}\right)+b\left(\frac{1}{x^2}\right)+c\left(\frac{1}{x^3}\right)\]

OpenStudy (roadjester):

Okay, so that's the "quote on quote" Dependence Equation, but how would I write it a a matrix to use the rref?

OpenStudy (anonymous):

You dont need to create a matrix at all for this problem.

OpenStudy (anonymous):

infact, this is an infinite dimensional vector space, so creating a matrix is pretty much out of the question >.<

OpenStudy (roadjester):

Well, mainly because that's how I learned how to do it, distribute the variables, in this case you used a, b, and c, and then write the constants in one row, the terms with powers of 1 in the next etc.

OpenStudy (anonymous):

right, that is correct when dealing with "regular" vector spaces. however, we have an abstract vector space before us. This calls for different techniques.

OpenStudy (roadjester):

what do you suggest?

OpenStudy (anonymous):

Here is an example. It will show you how to do this problem. I claim that:\[\frac{-3x^2-27x+2}{x^3}\]is in the span of those three functions. We can see this because:\[\frac{-3x^2-27x+2}{x^3}=\frac{-3x^2}{x^3}+\frac{-27x}{x^3}+\frac{3}{x^3}\]\[\frac{-3}{x}+\frac{-27}{x^2}+\frac{3}{x^3}=-3\cdot\frac{1}{x}-27\cdot\frac{1}{x^2}+3\cdot\frac{1}{x^3}\]So the function is a linear combination of 1/x, 1/x^2 and 1/x^3.

OpenStudy (roadjester):

uh, I get the denominator...but how did you get the numerator? I looks like you did partial fraction decomposition...

OpenStudy (anonymous):

All i did was divide it out. its like saying:\[\frac{5+4}{9}=\frac{5}{9}+\frac{4}{9}\]

OpenStudy (roadjester):

ah, in that case, 1) you're a fast typer to be able to type the fraction operator (using that annoying equation editor) 2) you have a typo.

OpenStudy (anonymous):

oh oh oh, yes, that 3 should be a 2. my bad.

OpenStudy (roadjester):

Ok, so since f(x) is a linear combination of the three "vectors" in S, it is in the span of the set?

OpenStudy (anonymous):

that is correct. you could even answer the "regular" vector space questions in this manner. if i asked if (1,3,5) was in the span of {(1,1,1) ,(1,2,3)}, you could either create the matrix, rref, etc, or just note that (1,3,5) = -1(1,1,1)+2(1,2,3). Both methods are correct.

OpenStudy (roadjester):

would one be easier than another?

OpenStudy (anonymous):

usually getting a matrix and row reducing (rref) can be a pain if done by hand. but it always works and is kinda mindless; just follow the set procedure. The other method usually involves thinking for a little, and sometimes you can get lucky and notice the linear combination. But sometimes its not so clear and the linear combination might not be obvious. I usually stare at the problem for a minute trying to guess the linear combination, and if I cant, then I create a matrix and rref, etc.